# Haoyuan Li,  Mathias Funk, Aaqib Saeed

### v1

@@ 6 Disuccusion
Discovering Novel Algorithmic Combinations. Helmsman advances complex problem-solving by strategically integrating human intelligence where full automation is insufficient. While standard tasks (e.g., Q4–Q8) are handled autonomously, complex queries such as Q16 initiate a structured human-in-the-loop protocol. This process begins with collaborative plan refinement to resolve ambiguity and concludes with human verification of the final synthesized code. This verification step is crucial, as our agile automated simulations may not capture the nuanced evaluation criteria of specialized research environments.
@@
FedAvg (McMahan et al., 2017) 16.38 0.75 7.83 0.76
FedProx (Li et al., 2020a) 16.71 0.76 8.19 0.75
FedEWC (Lee et al., 2017) 16.06 0.68 10.07 0.62
FedWeIT (Yoon et al., 2021) 28.56 0.49 20.48 0.45
FedLwF (Li & Hoiem, 2017) 30.94 0.42 21.74 0.46
TARGET (Zhang et al., 2023) 34.89 0.24 25.65 0.27
@@
This deliberate integration of human expertise is critical to our system’s success. In an extensive evaluation on Q16 (Table 5), the solution synthesized by Helmsman not only surpassed all competing Continual Learning methods but also demonstrated superior mitigation of catastrophic forgetting. This advantage is attributable to a novel combinatorial strategy that integrates client-side experience replay with global model distillation. This specific combination enhances knowledge sharing across continual learning tasks, enabling our hybrid approach to yield robust, deployment-ready solutions that overcome the limitations of standard research-focused methods.
Figure 4: Average evaluation attempts for Helmsman across 16 tasks with varying levels of human intervention. All results are averaged over 3 runs.
Computational Cost and the Limits of Autonomy. The primary limitation of our method is the computational cost associated with its iterative code generation and sandboxed federated simulation. To manage the intensive computation and communication inherent to this process, we restricted the maximum number of self-correction attempts during evaluation. Our results (Figure 4) show that Helmsman achieved full automation on 62.5% of benchmark tasks. The remaining tasks, however, require a degree of human intervention for successful completion. Therefore, the autonomy of our approach is currently bounded by the computational budget allocated for discovery, particularly for the most complex scientific problems.
@@ Appendix A Appendix
A.1 AgentFL-Bench Benchmark for Agentic System
@@
This section introduces details of AgentFL-Bench, a comprehensive benchmark for assessing the capabilities of agentic systems in the domain of federated learning research. The benchmark consists of 16 unique tasks, detailed in Tables 6, 7, and 8. These tasks are structured according to the query template in Table 1 and are organized into five key research domains, each characterized by a distinct set of challenges. Task queries are further categorized by their specific problem space using symbols, and essential information within each query is highlighted in green.
@@
A.2 Experimental Setup for Agentic Federated Learning Benchmark Evaluation
@@
A.3 Agent Configuration for Planning Stage
A.3.1 Prompt for Planning Agent
@@
A.3.2 Prompt for Reflection Agent
@@
A.4 Agent Configuration for Modular Coding Stage
A.4.1 Prompt for Supervision Agent
@@
A.4.2 Task Module Coding Group
@@
A.4.3 Strategy Module Coding Group
@@
A.4.4 Client Module Coding Group
@@
A.4.5 Server Module Coding Group
@@
A.4.6 Orchestrator Coding Group
@@
A.5 Agent Configuration for Evaluation Stage
A.5.1 Prompt for Evaluator Agent
@@
A.5.2 Prompt for Debugger Agent

### v2

@@ 6 Disuccusion
Discovering Novel Algorithmic Combinations. Helmsman advances complex problem-solving by strategically integrating human intelligence if available. While Helmsman can achieve 100% full automation (see Appendix A.1), complex queries such as Q16 can benefit from this human expertise via collaborative plan refinement to resolve ambiguity. As illustrated in Appendix A.3, our agile automated simulations can confront malicious or unaligned input queries.
@@
FedAvg (fedavg) 16.38 0.75 7.83 0.76
FedProx (fedprox) 16.71 0.76 8.19 0.75
FedEWC (fedewc) 16.06 0.68 10.07 0.62
FedWeIT (fedweit) 28.56 0.49 20.48 0.45
FedLwF (fedlwf) 30.94 0.42 21.74 0.46
TARGET (target) 34.89 0.24 25.65 0.27
@@
In the extensive evaluation on Q16 (Table 5), the solution synthesized by Helmsman surpassed all competing continual learning methods and exhibited markedly reduced catastrophic forgetting. This improvement stems from a novel combinatorial strategy that integrates client-side experience replay with global model distillation. By enhancing knowledge transfer across sequential tasks, this hybrid mechanism enables Helmsman to produce robust, deployment-ready solutions that overcome the limitations of conventional research-oriented approaches.
Evaluation on Helmsman System Components. To evaluate the contribution of each system component in Helmsman, we conduct an ablation study. As discussed in Section 3, Helmsman consists of three critical components. For a comprehensive comparison, we consider six configurations. We select seven representative query tasks from each research domain in AgentFL-Bench. As shown in Table 6, the single ReAct agent fails on all tasks. Although Configurations 2 and 3 solve a subset of queries, they fail on most, indicating that all system components are required for Helmsman to be fully functional. In contrast, the full Helmsman system achieves a 100% success rate. Notably, while removing the dual-layer verification reduces average API cost, it causes failures on all tasks, underscoring the essential role of verification in ensuring system robustness and stability.
[…]
Single ReAct Agent fail fail fail fail fail fail fail 0% 1.75
Single ReAct (w/ Dual Verification) fail fail success fail fail fail fail 14.29% 1.28
Helmsman (w/o Collab. Coding) fail fail fail success success fail fail 28.57% 2.11
Helmsman (w/o Dual Verification) fail fail fail fail fail fail fail 0% 0.88
Helmsman (Full System w/o HITL) success success success success success success success 100% 1.14
Helmsman (Full System) success success success success success success success 100% 0.98
@@ Appendix A Appendix
A.1 Evaluation Against State-of-the-Art Code Synthesis Methods
To further validate Helmsman’s effectiveness, we compare it against two state-of-the-art code-synthesis pipelines: Codex (GPT-5.1-codex) and Claude Code (Claude-Sonnet-4.5). We conduct additional experiments across all tasks in AgentFL-Bench using these baselines. As shown in Tables 7-11, Helmsman (with either GPT-5.1 or Claude-Sonnet-4.5) achieves a 100% success rate across all research queries, substantially outperforming Codex (37.5%) and Claude Code (43.75%). [^1] The configuration “Helmsman (Claude-Sonnet-4.0)” refers to earlier experiments conducted prior to the release of Sonnet 4.5.
Furthermore, Helmsman exhibits significantly lower API costs, benefiting from efficient agentic coordination throughout the research pipeline and thus surpassing existing code-automation approaches in cost efficiency. An additional observation is that GPT-5.1 serves as the most economical LLM backend for Helmsman, owing to its lower token consumption.
Table 7: Comprehensive performance comparison across federated learning tasks (Part a: Q1-Q3). For each task, we report Cost (\$, lower the better), Total Tokens (in thousands, lower the better), Walltime (seconds, lower the better), and Outcome. Best results per metric are in bold.
Q1 Q2 Q3
Method Cost (\$) Token Walltime (s) Outcome Cost (\$) Token Walltime (s) Outcome Cost (\$) Token Walltime (s) Outcome
Codex (GPT-5.1-Codex) 1.12 2,639k 834 success 1.64 4,339k 1,157 fail 1.12 2,683k 796 success
Claude Code (Claude Sonnet 4.5) 1.16 887k 634 fail 1.08 1,329k 1,617 success 1.02 969k 1,254 success
Helmsman (Claude Sonnet 4.0) 1.64 287k 841 success 2.02 372k 1,057 success 1.88 343k 773 success
Helmsman (Claude Sonnet 4.5) 1.53 221k 1,274 success 0.83 135k 1,154 success 0.91 164k 695 success
Helmsman (GPT-5.1) 0.64 213k 763 success 0.51 111k 587 success 0.62 190k 674 success
Table 8: Comprehensive performance comparison across federated learning tasks (Part b: Q4-Q6). For each task, we report Cost (\$, lower the better), Total Tokens (in thousands, lower the better), Walltime (seconds, lower the better), and Outcome. Best results per metric are in bold.
Q4 Q5 Q6
Method Cost (\$) Token Walltime (s) Outcome Cost (\$) Token Walltime (s) Outcome Cost (\$) Token Walltime (s) Outcome
Codex (GPT-5.1-Codex) 1.02 2,777k 743 fail 0.73 1,429k 618 success 1.43 3,939k 1,322 fail
Claude Code (Claude Sonnet 4.5) 1.69 2,245k 1,179 fail 2.38 3,165k 928 success 1.31 1,634k 1,436 fail
Helmsman (Claude Sonnet 4.0) 2.43 448k 914 success 1.39 253k 592 success 2.26 413k 865 success
Helmsman (Claude Sonnet 4.5) 0.68 134k 895 success 0.89 146k 469 success 1.66 200k 948 success
Helmsman (GPT-5.1) 0.72 199k 851 success 0.56 167k 592 success 0.80 295k 754 success
Table 9: Comprehensive performance comparison across federated learning tasks (Part c: Q7-Q9). For each task, we report Cost (\$, lower the better), Total Tokens (in thousands, lower the better), Walltime (seconds, lower the better), and Outcome. Best results per metric are in bold.
Q7 Q8 Q9
Method Cost (\$) Token Walltime (s) Outcome Cost (\$) Token Walltime (s) Outcome Cost (\$) Token Walltime (s) Outcome
Codex (GPT-5.1-Codex) 0.88 2,286k 867 fail 0.76 1,741k 968 fail 1.21 3,227k 1,405 success
Claude Code (Claude Sonnet 4.5) 1.85 1,484k 1,341 fail 1.87 3,273k 1,722 fail 2.61 3,002k 1,435 success
Helmsman (Claude Sonnet 4.0) 1.52 278k 923 success 1.13 207k 605 success 2.75 479k 781 success
Helmsman (Claude Sonnet 4.5) 0.94 362k 823 success 1.79 256k 774 success 1.20 209k 1,054 success
Helmsman (GPT-5.1) 0.64 324k 927 success 0.44 135k 896 success 0.51 150k 463 success
Table 10: Comprehensive performance comparison across federated learning tasks (Part d: Q10-Q13). For each task, we report Cost (\$, lower the better), Total Tokens (in thousands, lower the better), Walltime (seconds, lower the better), and Outcome. Best results per metric are in bold.
Q10 Q11 Q12 Q13
Method Cost (\$) Token Walltime (s) Outcome Cost (\$) Token Walltime (s) Outcome Cost (\$) Token Walltime (s) Outcome Cost (\$) Token Walltime (s) Outcome
Codex (GPT-5.1-Codex) 0.57 1,309k 726 success 0.97 2,453k 842 fail 0.62 1,538k 694 fail 0.70 1,498k 643 success
Claude Code (Claude Sonnet 4.5) 1.36 1,437k 1,171 success 2.54 2,714k 1,343 fail 0.93 1,042k 1,016 fail 2.44 5,862k 1,084 success
Helmsman (Claude Sonnet 4.0) 1.36 249k 819 success 1.94 357k 1,134 success 2.06 378k 1,066 success 1.98 364k 897 success
Helmsman (Claude Sonnet 4.5) 0.80 135k 792 success 0.86 150k 627 success 1.06 196k 571 success 1.27 204k 714 success
Helmsman (GPT-5.1) 0.63 160k 515 success 0.40 128k 938 success 0.57 165k 742 success 0.48 142k 594 success
Table 11: Comprehensive performance comparison across federated learning tasks (Part e: Q14-Q16 and Overall Summary). For each task, we report Cost (\$, lower the better), Total Tokens (in thousands, lower the better), Walltime (seconds, lower the better), and Outcome. The Average column shows mean performance across all 16 tasks with Success Rate. Best results per metric are in bold.
Q14 Q15 Q16 Average
Method Cost (\$) Token Walltime (s) Outcome Cost (\$) Token Walltime (s) Outcome Cost (\$) Token Walltime (s) Outcome Cost (\$) Token Walltime (s) Suc. Rate (%)
Codex (GPT-5.1-Codex) 0.57 1,519k 712 fail 0.44 1,073k 948 fail 1.09 4,831k 1,267 fail 0.93 2,455k 909 37.50
Claude Code (Claude Sonnet 4.5) 1.05 1,023k 1,227 success 1.79 2,844k 1,443 fail 2.14 2,823k 1,655 fail 1.70 2,233k 1,218 43.75
Helmsman (Claude Sonnet 4.0) 2.70 494k 1,229 fail 2.61 476k 966 fail 3.77 677k 1,368 fail 2.09 380k 927 81.25
Helmsman (Claude Sonnet 4.5) 0.74 133k 896 success 0.63 110k 1,159 success 0.90 149k 973 success 1.04 195k 864 100
Helmsman (GPT-5.1) 0.61 167k 713 success 0.39 107k 767 success 0.63 176k 682 success 0.57 177k 716 100
A.2 Evaluation on Reproducibility of Helmsman Framework
We further conduct an ablation study to assess the stability and reproducibility of Helmsman across three independent runs. We compare the final FL system structures generated by Helmsman with those produced by a state-of-the-art code-generation baseline (Claude Code). As shown in Figure 4, Helmsman consistently produces identical FL system structures across all runs, whereas Claude Code generates divergent system designs. This demonstrates the stability of Helmsman’s modularized architecture. Moreover, the modular design enables the Helmsman-generated FL systems to be easily adapted to alternative FL strategies by simply replacing the strategy module.
During the coding stage, a supervisor agent functions as the research lead, decomposing high-level research plans into implementable modules for downstream coding groups. As illustrated in Fig. 1, these coding groups operate in an interdependent manner, reflecting the intrinsic coupling between components in FL systems (e.g., client and server modules must coordinate to satisfy protocol requirements). The orchestration stage then ensures global logical consistency across the entire system. After verification by the orchestrator, the synthesized FL system proceeds to the evaluation stage, where a dual-layer sandboxed simulation is executed to validate system functionality and readiness for deployment.
(a) Claude Code (Run 1)
(b) Claude Code (Run 2)
(c) Claude Code (Run 3)
(d) Helmsman (All Runs)
Figure 4: Code generation stability comparison on Q1 task across 3 independent runs. Claude Code (Claude-Sonnet-4.5) produces distinct folder structures and implementations in each run (a-c), demonstrating inconsistent code generation. In contrast, Helmsman maintains identical system structure across all runs (d), ensuring reproducibility and enabling plug-and-play modularity.
A.3 Ablation Study of Planning Stability in Helmsman
A.3.1 The planning workflow of Human-Helmsman with Different Types of Input Schema
In Table 1, we introduce the structured research query pattern used in AgentFL-Bench. We identify three prerequisite components for constructing a complete FL query schema: the problem statement, task description, and framework requirements. This consistency in input structure ensures fairness in our experiments and supports a reliable evaluation of Helmsman’s performance in FL system generation. However, in real-world setting, the user input query schema can be misaligned with the tasks in AgentFL-Bench. Therefore, introduce the human-in-the-loop (HITP) in Helmsman to improve the system’s robustness.
To justify this, we conducted an additional ablation study evaluating Helmsman’s behavior under different structured input schemas (e.g., paraphrased, incomplete, and out-of-schema queries). During the planning stage, the planning team, which comprises the planner and self-reflection agents, collaborates to mitigate the risk of erroneous or unstable research plans. As show in the below Table 12 for incomplete input schemas, where essential information is missing for downstream synthesis, the planning team provides actionable feedback to the user to request the necessary details, as shown in the tables below. For paraphrased queries, Helmsman remains capable of producing valid research plans as long as the core information (as present in Table 1) is preserved. For out-of-schema inputs, Helmsman routes the query back to the user for a “sanity check,” thereby safeguarding the system from flawed assumptions or misaligned research trajectories.
We further show the complete interaction of Helmsman-Human when confronted with different types of misaligned input queries during the planning workflow using research query 1 as the example. Specifically, Figure 5 for original input query, Figure 6 for query type of out-of-schema, Figure 7 for query type of incomplete schema, and Figure 8 for query type of paraphrased schema.
Table 12: Helmsman’s Response to Different Input Schemas for Query 1 (CIFAR-10-LT FL Task)
Input Type User Query Helmsman Workflow
Original I need to deploy a photo-sorting app on 15 smartphones. Each phone stores a highly imbalanced slice of long-tail distributed data (CIFAR-10-LT). Help me build a federated learning framework to train a MobileNet-V1 model, evaluating performance by top-1 test accuracy. Plan Construction $\rightarrow$
Human Verification $\rightarrow$
Transition to Coding Group
Paraphrased I want to train a MobileNet model on 15 mobile devices that have unbalanced CIFAR-10-LT data. Can you help me set up a federated system and measure the accuracy? Plan Construction $\rightarrow$
Human Verification $\rightarrow$
Transition to Coding Group
Incomplete I need federated learning for 15 phones with imbalanced data. Model is MobileNet. Detect Missing Info $\rightarrow$
Request User Input $\rightarrow$
Receive Input $\rightarrow$
Plan Construction $\rightarrow$
Human Verification $\rightarrow$
Transition to Coding Group
Out-of-Schema Hello world!!!@#\$#\$% Detect Schema Anomaly $\rightarrow$
Manual Sanity Check $\rightarrow$
User Correction $\rightarrow$
Plan Construction $\rightarrow$
Human Verification $\rightarrow$
Transition to Coding Group
A.3.2 Human-Helmsman Conversation with Original Input
HELMSMAN System Starting Planning Workflow
HELMSMAN Input Required Enter your federated learning research question:
USER I need to deploy a photo-sorting app on 15 smartphones. Each phone stores a highly imbalanced slice of long-tail distributed data (CIFAR-10-LT). Help me build a federated learning framework to train a MobileNet-V1 model, evaluating performance by top-1 test accuracy.
HELMSMAN Agent: Planning Planning Agent Team: Thinking and researching… → Calling tools: web_search, docs_search → Web Search Results: {’query’: ’...’, ’answer’: ’...’, ’images’: [...], ’results’: ’...’} → Docs Search Results: {...}
HELMSMAN Agent: Reflection Initial Plan: Self-reflection: The plan is comprehensive and actionable for a research project. It clearly defines the objective, identifies the specific challenges (Non-IID, long-tail data), proposes a concrete methodology. It does not require further information from the user to proceed. Plan is complete and sound, routing to human for final verification…
HELMSMAN Approval Required Approve this plan? (yes/no)
USER Yes
HELMSMAN Complete Workflow Complete!
Figure 5: The complete interaction of Helmsman-Human during planning workflow on research query task 1. The input query type follows the original schema in AgentFL-Bench.
A.3.3 Human-Helmsman Conversation with Out-of-Schema Input
HELMSMAN System Starting Planning Workflow
HELMSMAN Input Required Enter your federated learning research question:
USER Hello world!!!@#\$#\$%
HELMSMAN Agent: Planning Planning Agent: Thinking and researching… Self-reflection: INCOMPLETE: The agent is explicitly asking for more information from the user and has not yet generated any part of the research plan (summary, challenges, tasks, or technical setup). Plan is incomplete, routing to human feedback…
HELMSMAN Refinement Required Agent Analysis: PLAN INCOMPLETE - Missing Information: The agent is explicitly asking for more information from the user and has not yet generated any part of the research plan (summary, challenges, tasks, or technical setup). CURRENT AVAILABLE INFORMATION: No information retrieved yet Please provide additional information to help create a complete FL research plan. Awaiting for user input:
USER I need to …
HELMSMAN Agent: Reflection Initial Plan: Self-reflection: Plan is complete, routing to human decision…
HELMSMAN Approval Required Approve this plan? (yes/no)
USER Yes
HELMSMAN Complete Workflow Complete!
Figure 6: The complete interaction of Helmsman-Human during planning workflow on research query task 1. The input query type follows the Out-of-Schema setting in AgentFL-Bench.
A.3.4 Human-Helmsman Conversation with Incomplete Input
HELMSMAN System Starting Planning Workflow
HELMSMAN Input Required Enter your federated learning research question:
USER I need federated learning for 15 phones with imbalanced data. Model is MobileNet.
HELMSMAN Agent: Planning Planning Agent: Thinking and researching… → Calling tools: web_search, docs_search → Web Search Results: {’query’: ’...’, ’answer’: ’...’, ’images’: [...], ’results’: ’...’} → Docs Search Results: {...}
HELMSMAN Agent: Reflection Initial Plan: Self-reflection: INCOMPLETE: The plan is missing specific details for the dataset, the precise data partition strategy for imbalance, and the primary evaluation criteria, as explicitly requested by the agent in its final message. PLAN INCOMPLETE - Missing Information: The plan is missing specific details for the dataset, the precise data partition strategy for imbalance, and the primary evaluation criteria, as explicitly requested by the agent in its final message. Plan is incomplete, routing to human feedback…
HELMSMAN Refinement Required Provide additional information or requirements:
USER The dataset is …
HELMSMAN Agent: Reflection Initial Plan: Self-reflection: Plan is complete, routing to human decision…
HELMSMAN Approval Required Approve this plan? (yes/no)
USER Yes
HELMSMAN Complete Workflow Complete!
Figure 7: The complete interaction of Helmsman-Human during planning workflow on research query task 1. The input query type follows the Incomplete schema in AgentFL-Bench.
A.3.5 Human-Helmsman Conversation with Paraphrased Input
HELMSMAN System Starting Planning Workflow
HELMSMAN Input Required Enter your federated learning research question:
USER I want to train a MobileNet model on 15 mobile devices that have unbalanced CIFAR-10-LT data. Can you help me set up a federated system and measure the accuracy?
HELMSMAN Agent: Planning Planning Agent Team: Thinking and researching… → Calling tools: web_search, docs_search → Web Search Results: {’query’: ’...’, ’answer’: ’...’, ’images’: [...], ’results’: ’...’} → Docs Search Results: {...}
HELMSMAN Agent: Reflection Initial Plan: Self-reflection: The plan is comprehensive and actionable for a research project. It clearly defines the objective, identifies the specific challenges (Non-IID, long-tail data), proposes a concrete methodology. It does not require further information from the user to proceed. Plan is complete and sound, routing to human for final verification…
HELMSMAN Approval Required Approve this plan? (yes/no)
USER Yes
HELMSMAN Complete Workflow Complete!
Figure 8: The complete interaction of Helmsman-Human during planning workflow on research query task 1. The input query type follows the paraphrased schema in AgentFL-Bench.
A.4 AgentFL-Bench Benchmark for Agentic System
@@
This section introduces details of AgentFL-Bench, a comprehensive benchmark for assessing the capabilities of agentic systems in the domain of federated learning research. The benchmark consists of 16 unique tasks, detailed in Tables 13, 14, and 15. These tasks are structured according to the query template in Table 1 and are organized into five key research domains, each characterized by a distinct set of challenges. Task queries are further categorized by their specific problem space using symbols, and essential information within each query is highlighted in green.
@@
A.5 Experimental Setup for Agentic Federated Learning Benchmark Evaluation
@@
A.6 Agent Configuration for Planning Stage
A.6.1 Prompt for Planning Agent
@@
A.6.2 Prompt for Reflection Agent
@@
A.7 Agent Configuration for Modular Coding Stage
A.7.1 Prompt for Supervision Agent
@@
A.7.2 Task Module Coding Group
@@
A.7.3 Strategy Module Coding Group
@@
A.7.4 Client Module Coding Group
@@
A.7.5 Server Module Coding Group
@@
A.7.6 Orchestrator Coding Group
@@
A.8 Agent Configuration for Evaluation Stage
A.8.1 Prompt for Evaluator Agent
@@
A.8.2 Prompt for Debugger Agent